A normal distribution has μ = 80 and σ = 10. What is the probability of randomly selecting between the mean and a score of 90


1
Expert's answer
2021-06-03T15:21:12-0400

"P(80<X<90)=P(\frac{80-80}{10}<Z<\frac{90-80}{10})=P(0<Z<1)="

"=P(Z<1)-P(Z<0)=0.8413-0.5=0.3413."